Solve: x^2y^2(2ydx + xdy) - (5ydx + 7xdy)=0
1
Expert's answer
2020-09-10T19:23:14-0400

Solution. We write the equation as


"2x^2y^3dx+x^3y^2dy-5ydx-7xdy=0"

Let's select the total differential x^3y^3


"\frac{1}{3} d(x^3y^3)+x^2y^3dx-5ydx-7xdy=0."

Let's select the total differential -5xy


"\frac{1}{3} d(x^3y^3)-5d(xy)+x^2y^3dx-2xdy=0."

Consider the equation


"x^2y^3dx-2xdy=0"

get


"xdx=\frac{2dy}{y^3}."

As result


"\frac{x^2}{2}+\frac{1}{y^2}=C"

where C is constant.

Hence. the general solution of the equation is


"\frac{x^3y^3}{3}-5xy+\frac{x^2}{2}+\frac{1}{y^2}=c"

Answer.

"\frac{x^3y^3}{3}-5xy+\frac{x^2}{2}+\frac{1}{y^2}=c"